# DeDisCo at the DISRPT 2025 Shared Task: A System for Discourse Relation Classification

### v2

@@ 3 Data and Approach
The Discourse Relation Classification task of the DISRPT 2025 aims to classify discourse relations across a diverse set of languages and annotation frameworks. The relation classification task includes 38 of the joint task’s 39 corpora, spanning 16 languages and 6 different frameworks, with a unified set of 17 labels provided. For convenience, the language, corpora, and framework are part of the listed in appendix˜A (footnote 5). The majority of the data are annotated using Rhetorical Structure Theory (RST; Mann and Thompson, 1988) and the Penn Discourse Treebank framework (PDTB; Marcu and Wong, 2004). A smaller subset of the corpora is annotated using SDRT, discourse dependencies (DEP), eRST, and the ISO frameworks.
@@ 3.1 Features
As noted, the training data spans multiple annotation frameworks, corpora, and languages, which poses a challenge for generalization since relations defined under different schemes or languages are not always directly comparable. To address this diversity, we incorporate metadata into the input to help the model distinguish and generalize across datasets. We refer to these elements collectively as LCF features (Language, Corpus, Framework). For example, the dataset identifier zho.rst.gcdt indicates a Chinese corpus (zho) named GCDT Peng et al. (2022), annotated using the RST framework.
@@ 3.3 Pruned Qwen3-4B Decoder Only
We also experimented with mT5 Xue et al. (2021), a multilingual T5 variant pretrained on a Common Crawl-based corpus covering 101 languages. Specifically, we selected the mT5-XL variant, comprising 3.7 billion parameters, which comes closest to the shared task 4B limit. For our classification task, we used only the encoder and added a classification head. We explored two strategies for incorporating metadata and discourse features: (1) encoding them as special input tokens, and (2) using separate embedding layers concatenated with the encoder input.
@@ 4.1 Performance Comparison of Encoder-Only and Decoder-Only Model
Table 3: Accuracy of encoder-only and decoder-only models on the test sets of 38 corpora. We use corpus codes for simplicity (check footnote 5 in appendix˜A for language and framework information), with accuracy scores reported separately for the decoder-only and encoder-only models. Macro and micro averages are reported at the bottom.
@@
We also observed differences in how the two architectures respond to feature integration. LCF and DiscoDisco features consistently improved the encoder model, but in many cases degraded performance for the decoder. This suggests that encoder-only models can more effectively leverage categorical metadata and structural cues as additional signals, whereas decoder-only models are more sensitive to such injections. In contrast, extending the context window benefited the decoder but often harmed the encoder.
@@ 4.2 Decoder Model Ablation Tests
To assess the contribution of each additional features, we conduct an ablation study on the decoder-only model. In each experiment, one specific feature is removed from the input to evaluate its impact on performance. Detailed results are presented in appendix˜DTable 7).
We find that direction is the most influential feature. When direction information is removed, 32 out of the 38 corpora experience a drop in accuracy of more than 4%, and 12 corpora suffer a decrease of over 10%. By contrast, for tha.pdtb.tdtb, direction contributes only marginal gains, and for tur.pdtb.tedm, it even leads to a performance drop. This is likely because Turkish is a free-word-order left-branching language, where in most subordinate constructions the clause with the connective precedes the main clause, causing direction features based on linear order to introduce noise rather than useful signal Zeyrek et al. (2009).
@@ 6 Conclusion
Appendix A Corpora from DISRPT 2025 Shared Task
@@
Table 4: Collation of the data in DISRPT 2025 Shared task with the corresponding Language and Framework for reference. The frameworks include RST Mann and Thompson (1988), PDTB Marcu and Wong (2004), DEP Stede et al. (2016), SDRT Lascarides and Asher (2007), eRST Zeldes et al. (2025), and ISO 24617-8 Bunt and Prasad (2016); Tomaszewska et al. (2024). The languages are sorted based on the language code (ISO 639-3555https://www.iso.org/iso-639-language-code). The corpus fra.sdrt.summre marked with $\prime{}^{\prime}$ is not part of the relation classification task.
Appendix B Feature Utilization in Decoder and Encoder Models
Feature Group Decoder Encoder
LCF features + +
DiscoDisco features + +
Direction + +
Context (window size) +
Augmented dataset + +
DiscoDisco Feature Decoder Encoder
Genre +
Children
Discontinuous +
Is Sentence +
Length Ratio
Same Speaker + +
Document Length
Position +
Distance +
Lexical Overlap
Appendix C Experimental Setup of Decoder
Appendix D Ablation Test Results of Decoder

### v3

@@ 3 Data and Approach
The Discourse Relation Classification task of the DISRPT 2025 aims to classify discourse relations across a diverse set of languages and annotation frameworks. The relation classification task includes 38 of the joint task’s 39 corpora, spanning 16 languages and 6 different frameworks, with a unified set of 17 labels provided. For convenience, the language, corpora, and framework are part of the listed in appendix˜D (footnote 6). The majority of the data are annotated using Rhetorical Structure Theory (RST; Mann and Thompson, 1988) and the Penn Discourse Treebank framework (PDTB; Marcu and Wong, 2004). A smaller subset of the corpora is annotated using SDRT, discourse dependencies (DEP), eRST, and the ISO frameworks.
@@ 3.1 Features
As noted, the training data spans multiple annotation frameworks, corpora, and languages, which poses a challenge for generalization since relations defined under different schemes or languages are not always directly comparable. To address this diversity, we incorporate metadata into the input to help the model distinguish and generalize across datasets. We refer to these elements collectively as LCF features (Language, Corpus, Framework). For example, the dataset identifier zho.rst.gcdt indicates a Chinese corpus (zho) named GCDT Peng et al. (2022b), annotated using the RST framework.
@@ 3.3 Pruned Qwen3-4B Decoder Only
We also experimented with mT5 Xue et al. (2021), a multilingual T5 variant pretrained on a Common Crawl-based corpus covering 101 languages. Specifically, we selected the mT5‑XL variant, comprising 3.7 billion parameters, which comes closest to the shared task 4B limit. For our classification task, we used only the encoder and added a classification head. We explored two strategies for incorporating metadata and discourse features: (1) encoding them as special input tokens, and (2) using separate embedding layers concatenated with the encoder input.
@@ 4.1 Performance Comparison of Encoder-Only and Decoder-Only Model
Table 3: Accuracy of encoder-only and decoder-only models on the test sets of 38 corpora. We use corpus codes for simplicity (check footnote 6 in appendix˜D for language and framework information), with accuracy scores reported separately for the decoder-only and encoder-only models. Macro and micro averages are reported at the bottom.
@@
We also observed differences in how the two architectures respond to feature integration. LCF and DiscoDisco features consistently improved the encoder model, but in many cases degraded performance for the decoder. This suggests that encoder‑only models can more effectively leverage categorical metadata and structural cues as additional signals, whereas decoder‑only models are more sensitive to such injections. In contrast, extending the context window benefited the decoder but often harmed the encoder.
@@ 4.2 Decoder Model Ablation Tests
To assess the contribution of each additional features, we conduct an ablation study on the decoder-only model. In each experiment, one specific feature is removed from the input to evaluate its impact on performance. Detailed results are presented in appendix˜ETable 7).
We find that direction is the most influential feature. When direction information is removed, 32 out of the 38 corpora experience a drop in accuracy of more than 4%, and 12 corpora suffer a decrease of over 10%. By contrast, for tha.pdtb.tdtb, direction contributes only marginal gains, and for tur.pdtb.tedm, it even leads to a performance drop. This is likely because Turkish is a free-word-order left‑branching language, where in most subordinate constructions the clause with the connective precedes the main clause, causing direction features based on linear order to introduce noise rather than useful signal Zeyrek et al. (2009).
@@ 6 Conclusion
Appendix A Data
We train and evaluate our models using all the datasets provided by the shared task organizers. [^4] https://github.com/disrpt/sharedtask2025 In total, the benchmark is composed of $39$ datasets, covering $13$ languages and $6$ frameworks. These datasets were obtained from the following corpora: the Czech RST Discourse Treebank 1.0 Poláková et al. (2023), the Potsdam Commentary Corpus Stede and Neumann (2014); Bourgonje and Stede (2020), the COVID-19 Discourse Dependency Treebank Nishida and Matsumoto (2022), the Discourse Dependency TreeBank for Scientific Abstracts Yang and Li (2018); Yi et al. (2021); Cheng and Li (2019), the Genre Tests for Linguistic Evaluation corpus Aoyama et al. (2023), the Georgetown University Multilayer corpus Zeldes (2017), the RST Discourse Treebank Carlson et al. (2001), the Science, Technology, and Society corpus Potter (2008), the University of Potsdam Multilayer UNSC Corpus Zaczynska and Stede (2024), the Minecraft Structured Dialogue Corpus Thompson et al. (2024), the Strategic Conversations corpus Asher et al. (2016), the Basque RST Treebank Iruskieta et al. (2013), the Persian RST Corpus Shahmohammadi et al. (2021), the ANNOtation DIScursive corpus Afantenos et al. (2012), the SUMM-RE corpus Hunter et al. (2024); Prévot et al. (2025), the Dutch Discourse Treebank Redeker et al. (2012), the Polish Discourse Corpus Ogrodniczuk et al. (2024); Calzolari et al. (2024), the Cross-document Structure Theory News Corpus Cardoso et al. (2011), the Russian RST Treebank Toldova et al. (2017), the RST Spanish Treebank da Cunha et al. (2011), the RST Spanish-Chinese Treebank Cao et al. (2018), the Georgetown Chinese Discourse Treebank Peng et al. (2022b, a), the DiscoNaija corpus Scholman et al. (2025), the Penn Discourse Treebank Prasad et al. (2014); Webber et al. (2019), the TED-Multilingual Discourse Bank (English) Zeyrek et al. (2018, 2019), the LUNA Corpus Discourse Data Set Tonelli et al. (2010); Riccardi et al. (2016), the Portuguese Discourse Bank Mendes and Lejeune (2022); Généreux et al. (2012), the Thai Discourse Treebank Prasertsom et al. (2024), the Turkish Discourse Bank Zeyrek and Webber (2008); Zeyrek and Kurfalı (2017), and the Chinese Discourse Treebank Zhou et al. (2014).
Appendix B Feature Utilization in Decoder and Encoder Models
Feature Group Decoder Encoder
LCF features + +
DiscoDisco features + +
Direction + +
Context (window size) +
Augmented dataset + +
DiscoDisco Feature Decoder Encoder
Genre +
Children
Discontinuous +
Is Sentence +
Length Ratio
Same Speaker + +
Document Length
Position +
Distance +
Lexical Overlap
Appendix C Experimental Setup of Decoder
Appendix D Corpora from DISRPT 2025 Shared Task
@@
Table 6: Collation of the data in DISRPT 2025 Shared task with the corresponding Language and Framework for reference. The frameworks include RST Mann and Thompson (1988), PDTB Marcu and Wong (2004), DEP Stede et al. (2016), SDRT Lascarides and Asher (2007), eRST Zeldes et al. (2025), and ISO 24617-8 Bunt and Prasad (2016); Tomaszewska et al. (2024). The languages are sorted based on the language code (ISO 639-3666https://www.iso.org/iso-639-language-code). The corpus fra.sdrt.summre marked with $\prime{}^{\prime}$ is not part of the relation classification task.
Appendix E Ablation Test Results of Decoder